OpenStudy (anonymous):

subtract and simplify: (9x/3x-4) - (12/3x-4)

OpenStudy (anonymous):

\[ \frac{9x}{3x-4}-\frac{12}{3x-4}=\frac{9x-12}{3x-4}\]